40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 23 15:26:54 2018, Jon, Update, IMC, MC/PMC trouble 
    Reply  Thu Aug 23 16:05:24 2018, Koji, Update, IMC, MC/PMC trouble 
       Reply  Thu Aug 23 16:10:13 2018, not Koji, Update, IMC, MC/PMC trouble 
Message ID: 14179     Entry time: Thu Aug 23 15:26:54 2018     Reply to this: 14180
Author: Jon 
Type: Update 
Category: IMC 
Subject: MC/PMC trouble 

I tried unsuccessfully to relock the MC this afternoon.

I came in to find it in a trouble state with a huge amount of noise on C1:PSL-FSS_PCDRIVE visible on the projector monitor. Light was reaching the MC but it was unable to lock.

  • I checked the status of the fast machines on the CDS>FE STATUS page. All up.
  • Then I checked the slow machine status. c1iscaux and c1psl were both down. I manually reset both machines. The large noise visible on C1:PSL-FSS_PCDRIVE disappeared.
  • After the reset, light was no longer reaching the MC, which I take to mean the PMC was not locked. On the PSL>PMC page, I blanked the control signal, reenabled it, and attempted to relock by adjusting the servo gain as Gautam had showed me before. The PMC locks were unstable, with each one lasting only a second or so.
  • Next I tried restoring the burt states for c1iscaux and c1psl from a snapshot taken earlier today, before the machine reboots. That did not solve the problem either.
ELOG V3.1.3-